# A Fast AI Surrogate for Coastal Ocean Circulation Models

### v1

@@ I Introduction
The emergence of Graphics Processing Units (GPUs) has enabled massive parallelism in computing. Several new national supercomputers, including Summit, now rely predominantly on GPU-based computation. However, directly porting ROMS MPI codes to the CUDA GPU architecture is non-trivial. The spatial domain can be a non-uniform mesh with varying cell densities. Converting ROMS codes from MPI to GPUs requires additional design and optimization in memory coalescing and thread synchronization. Furthermore, ROMS parallelism is based on MPI, which cannot be directly ported to GPUs. There are several efforts to support ROMS on GPUs [9, 10]. However, the improvement of these methods is limited, and they fail to exploit the capabilities of modern GPU hardware.
@@
To address these challenges, this paper introduces an AI surrogate based on a 4D Swin Transformer for coastal circulation modeling in both hindcasts and forecasts (up to 12 days). Coastal circulation is primarily driven by astronomic tides, wind stress, atmospheric pressure gradients, water density gradients, and freshwater inputs, all of which shape the flow and mixing patterns along coastlines. In this study, to simplify the problem and reduce the complexity of the processes driving the coastal circulation, we focus on characterizing the water level and the flow associated with tidal propagation. We develop a fully GPU-accelerated workflow, optimizing the model training and inference pipeline on NVIDIA DGX A100 GPUs. Our approach not only accelerates simulations but also incorporates a physics-based constraint (the conservation of mass) to detect and correct inaccurate results, ensuring reliability while minimizing manual intervention. Our experiments demonstrate that the AI surrogate achieves up to 450$\times$ speedup over traditional ROMS simulations while maintaining high-quality simulation results. This paper makes the following contributions:
@@ II-A AI Surrogate for physical models
Existing machine learning techniques for numerical simulations include physics-informed neural networks (PINNs) and neural operator learning. A PINN trains a neural network to solve a partial different equation (PDE) through the automatic differentiation [25]. PINNs have been tested in simulating coastal flood and wave [26, 27], but their training is challenging due to the complex physics-informed regularization terms [28] and the huge computational cost for solving the PDE [29]. Neural operator learning aims to train a neural network surrogate as the solver of a family of PDE instances [30]. The surrogate takes the initial or boundary conditions and predicts the solution function. Existing surrogate models include deep convolutional neural networks [31], graph neural operators [32, 33], Fourier neural operators [34, 35], DeepONet [30], NodeFormer [36, 37] and vision transformers [38, 21]. Some simple methods have been tested for individual coastal problems (e.g., storm surges) [39, 40]. There are also methods for irregular spatial points or unstructured meshes in continuous space through implicit neural representation [41, 42, 43] or geometric deformation [44], or fixed-graph transformation [45], but their neural networks are limited by only taking each sample’s spatial coordinates without explicitly capturing spatial structure. There also exists a rich literature that particularly focuses on machine learning for continuous fluid dynamics (CFD) and turbulence [46], e.g., accelerating direct numerical simulations [47], improving turbulence closure modeling [48], and enhancing reduced-order models [49, 50]. However, existing research mostly focuses on idealized settings, with few works addressing real-world complex coastal circulation in estuaries.
@@ III-A Overall Workflow
The overall workflow is shown in Figure 1. Historical coastal simulations from ROMS are pre-processed and used to train an AI surrogate. Once trained, the AI surrogate can make forecasts by taking the initial condition of physical variables at time $t_{0}$ and the boundary conditions of those variables from time $t_{1}$ to time $t_{T}$ as inputs. The surrogate predicts the interior values of those variables at time $t_{1}$ to $t_{T}$ as outputs. Some variables, such as free surface elevation, are 2D, while others, such as circulation velocities, are 3D. A physics-based verification module checks if the surrogate’s forecast satisfies the physical constraints, such as the conservation of mass. If the forecasts meet these requirements, the forecasting continues with the surrogate generating the next episode of snapshots. Otherwise, the pipeline temporarily switches back to the original ROMS model to run the forecast.
For long-term forecasting, we employ a dual-model approach, i.e., training two models with the same architecture but at different temporal resolutions. Specifically, one model makes forecasts over 12 days with 24 snapshots (one snapshot every 12 hours), while the other model makes forecasts over 12 hours with 24 snapshots (one snapshot every half an hour). To generate 12-day forecasts with half-hour intervals (576 snapshots), we combine the two models, i.e., first running the coarse model to generate a 12-day forecast with a snapshot every half day and then using each coarse snapshot as the initial condition to generate 24 half-hour snapshots via the second model.
@@ III-C AI surrogate architecture
Our AI surrogate model is built upon the 4D Swin Transformer model [54, 55, 56]. It has an encoder-decoder architecture, as shown in Figure 2. The model starts by taking the initial conditions at time $t=0$ and boundary conditions from $t=1$ to $T$ across all simulation variables (both 3D variables $u,v,w$ and 2D variable $\zeta$ on the mesh surface). To efficiently process this high-dimensional data with millions of mesh cells, we first employ patch partitioning and embedding for 3D and 2D variables. Patch partitioning divides the 3D/2D spatial domain into smaller blocks, effectively reducing the spatial resolution. Patch embedding transforms these blocks into the same latent representation space, facilitating the concatenation of a 2D variable (free surface elevation $\zeta$) and 3D variables (current velocity $u,v,w$) along the depth dimension. After considering an additional temporal dimension, the 4D data was then fed into a 4D Swin Transformer encoder. It leverages the 4D spatiotemporal attention mechanism to learn complex dependency patterns across both space and time with advanced window self-attention and shifted-window self-attention layers for computational efficiency, simultaneously downsampling the data with patch merging operations to capture the multi-scale information. The decoder employs multiple layers of transposed convolutions, with batch normalization [57] and Gaussian Error Linear Units (GELU) [58] activations to gradually upsample the encoded data, restoring it to its original resolution. A skip connection operation similar to the U-Net architecture [59, 60] is introduced to fuse fine-grained information directly from the corresponding encoder layers. Finally, the model splits the 3D and 2D variables after upsampling and recovers them to the original spatial and temporal shape.
@@ III-E Inference and verification
Verification to ensure adherence to physical laws. At the core of our verification strategy is the mass conservation law of water. To satisfy this law, the divergence of circulation velocities into a water column (water inflow) should be equal to the time derivative of free surface elevation (increase of volume). For a horizontal domain $\Omega$ of contour $\Gamma$, the mass conservation law can be expressed as Equation 4,
@@ IV-C Case Study
(a) Horizontal current velocity $u$ (east-west)
(b) Horizontal current velocity $v$ (north-south)
(c) Free surface elevation $\zeta$
(a) Location 1 ($26.35\degree N,82.06\degree W$)
(b) Location 2 ($26.72\degree N,82.25\degree W$)
(c) Location 3 ($26.84\degree N,82.13\degree W$)
@@ IV-E Sensitivity Analysis
TABLE IV: Sensitivity analysis of the patch size.
Patch Size # of Parameters [M] Time/Instance [s] MAE RMSE
$u$ [m/s] $v$ [m/s] $w$ [m/s] $\zeta$ [m] $u$ [m/s] $v$ [m/s] $w$ [m/s] $\zeta$ [m]

### v2

# Accelerate Coastal Ocean Circulation Model with AI Surrogate

@@ I Introduction
The emergence of Graphics Processing Units (GPUs) has enabled massive parallelism in computing. Several national supercomputers, including Summit, now rely predominantly on GPU-based computation. However, directly porting ROMS MPI codes to the CUDA GPU architecture is non-trivial. The spatial domain can be a non-uniform mesh with varying cell densities. Converting ROMS codes from MPI to GPUs requires additional design and optimization in memory coalescing and thread synchronization. Furthermore, ROMS parallelism is based on MPI, which cannot be directly ported to GPUs. There are several efforts to support ROMS on GPUs [9, 10, 11]. However, the improvement of these methods is limited, and they fail to exploit the full capabilities of modern GPU hardware.
@@
To address these challenges, this paper introduces an AI surrogate based on a 4D Swin Transformer for coastal circulation modeling in both hindcasts (predicting past variables from historical data) and forecasts (predicting future variables) up to 12 days. Coastal circulation is primarily driven by astronomic tides, wind stress, atmospheric pressure gradients, water density gradients, and freshwater inputs, all of which shape the flow and mixing patterns along coastlines. In this study, to simplify the problem and reduce the complexity of the processes driving the coastal circulation, we focus on characterizing the water level and the flow associated with tidal propagation. We develop a fully GPU-accelerated workflow, optimizing the model training and inference pipeline on NVIDIA DGX A100 GPUs. Our approach not only accelerates simulations but also incorporates a physics-based constraint (the conservation of mass) to detect and correct inaccurate results, ensuring reliability while minimizing manual intervention. Our experiments demonstrate that the AI surrogate achieves up to 450$\times$ speedup over traditional ROMS simulations while maintaining high-quality simulation results. The code and data are available at https://github.com/spatialdatasciencegroup/CoastalOceanAISurrogate. This paper makes the following contributions:
@@ II-A AI Surrogate for physical models
Existing machine learning techniques for numerical simulations include physics-informed neural networks (PINNs) and neural operator learning. A PINN trains a neural network to solve a partial different equation (PDE) through the automatic differentiation [26]. PINNs have been tested in simulating coastal flood and wave [27, 28], but their training is challenging due to the complex physics-informed regularization terms [29] and the huge computational cost for solving the PDE [30]. Neural operator learning aims to train a neural network surrogate as the solver of a family of PDE instances [31]. The surrogate takes the initial or boundary conditions and predicts the solution function. Existing surrogate models include deep convolutional neural networks [32], graph neural operators [33, 34], Fourier neural operators [35, 36], DeepONet [31], NodeFormer [37, 38] and vision transformers [39, 22]. Some simple methods have been tested for individual coastal problems (e.g., storm surges) [40, 41]. There are also methods for irregular spatial points or unstructured meshes in continuous space through implicit neural representation [42, 43, 44] or geometric deformation [45], or fixed-graph transformation [46], but their neural networks are limited by only taking each sample’s spatial coordinates without explicitly capturing spatial structure. There also exists a rich literature that particularly focuses on machine learning for computational fluid dynamics (CFD) and turbulence [47], e.g., accelerating direct numerical simulations [48], improving turbulence closure modeling [49], and enhancing reduced-order models [50, 51]. However, existing research mostly focuses on idealized settings, with few works addressing real-world complex coastal circulation in estuaries.
@@ III-A Overall Workflow
The overall workflow is shown in Figure 1. Historical coastal simulations from ROMS are pre-processed and used to train an AI surrogate. Once trained, the AI surrogate can make forecasts by taking the initial condition of physical variables at time $t_{0}$ and the boundary conditions of those variables from time $t_{1}$ to time $t_{T}$ as inputs. For example, to forecast 24 snapshots (one snapshot every half hour) of variables inside a 3D mesh from 12:00 p.m. on March 26, 2012, the initial condition is the state of variables on the entire 3D mesh at a previous snapshot (11:30 a.m. on March 26, 2012). Boundary conditions include variables at the lateral boundary of the 3D mesh in the 24 snapshots to be forecasted, spanning up to 12:00 a.m. on March 27, 2012. The surrogate predicts the interior values of those variables at time $t_{1}$ to $t_{T}$ as outputs. Some variables, such as free surface elevation, are 2D, while others, such as circulation velocities, are 3D. A physics-based verification module checks if the surrogate’s forecast satisfies the physical constraints, such as the conservation of mass. If the forecasts meet these requirements, the forecasting continues with the surrogate generating the next episode of snapshots. Otherwise, the pipeline temporarily switches back to the original ROMS model to run the forecast.
For long-term forecasting, we employ a dual-model approach, i.e., training two models with the same architecture but at different temporal resolutions. Specifically, one model makes forecasts over 12 days with 24 snapshots (one snapshot every 12 hours), while the other model makes forecasts over 12 hours with 24 snapshots (one snapshot every half an hour). To generate 12-day forecasts with half-hour intervals (576 snapshots), we combine the two models, i.e., first running the coarse model to generate a 12-day forecast with a snapshot every half day and then using each coarse snapshot as the initial condition to generate 24 half-hour snapshots via the second model. We set 12 days as the prediction limit according to oceanographers for practical use. The 12-day prediction limit is set based on practical considerations from oceanographers, as boundary condition uncertainties increase over time, making longer forecasts increasingly unreliable, even for ROMS.
@@ III-C AI surrogate architecture
Our AI surrogate model is built upon the 4D Swin Transformer model [55, 56, 57]. It has an encoder-decoder architecture, as shown in Figure 2. The model starts by taking the initial conditions at time $t=0$ and boundary conditions from $t=1$ to $T$ across all simulation variables (both 3D variables $u,v,w$ and 2D variable $\zeta$ on the mesh surface). To efficiently process this high-dimensional data with millions of mesh cells, we first employ patch partitioning and embedding [58] for 3D and 2D variables. Patch partitioning divides the 3D/2D spatial domain into smaller blocks, effectively reducing the spatial resolution. Patch embedding transforms these blocks into the same latent representation space, facilitating the concatenation of a 2D variable (free surface elevation $\zeta$) and 3D variables (current velocity $u,v,w$) along the depth dimension. After considering an additional temporal dimension, the 4D data was then fed into a 4D Swin Transformer encoder. It leverages the 4D spatiotemporal attention mechanism to learn complex dependency patterns across both space and time with advanced window self-attention and shifted-window self-attention layers for computational efficiency, simultaneously downsampling the data with patch merging operations to capture the multi-scale information. The decoder employs multiple layers of transposed convolutions, with batch normalization [59] and Gaussian Error Linear Units (GELU) [60] activations to gradually upsample the encoded data, restoring it to its original resolution. A skip connection operation similar to the U-Net architecture [61, 62] is introduced to fuse fine-grained information directly from the corresponding encoder layers. Finally, the model splits the 3D and 2D variables after upsampling and recovers them to the original spatial and temporal shape.
@@ III-E Inference and verification
Verification to ensure adherence to physical laws. At the core of our verification strategy is the mass conservation law of water, as it is straightforward and reliably tests one important aspect of the AI model’s physical consistency. To satisfy this law, the divergence of circulation velocities into a water column (water inflow) should be equal to the time derivative of free surface elevation (increase of volume). For a horizontal domain $\Omega$ of contour $\Gamma$, the mass conservation law can be expressed as Equation 4,
@@ IV Evaluation
(a) Horizontal current velocity $u$ (east-west)
(b) Horizontal current velocity $v$ (north-south)
(c) Free surface elevation $\zeta$
@@ IV-A System hardware platform
(a) Selected Locations
(b) Location 1 ($26.35\degree N,82.06\degree W$)
(c) Location 2 ($26.49\degree N,82.03\degree W$)
(d) Location 3 ($26.72\degree N,82.24\degree W$)
@@ IV-C Case Study
TABLE IV: Sensitivity analysis of the patch size.
Patch Size # of Parameters [M] Time/Instance [s] MAE RMSE
$u$ [m/s] $v$ [m/s] $w$ [m/s] $\zeta$ [m] $u$ [m/s] $v$ [m/s] $w$ [m/s] $\zeta$ [m]
@@ V Conclusion and Future Work
Acknowledgments
This material is based upon work supported by the National Science Foundation (NSF) under Grant No. IIS-2147908, IIS-2207072, OAC-2152085, OAC-2402946, OAC-2410884, and OAC-2402947.